# DEFAME: Dynamic Evidence-based FAct-checking with Multimodal Experts

### v2

@@ E.3 Summarize Prompt Template
E.4 Develop Prompt Template
Instructions You just retrieved new Evidence. Now, analyze the Claim’s veracity using the evidence. Always adhere to the following rules: Focus on developing new insights. Do not repeat larger parts from the Record. Do not restate the Claim. Write down your thoughts step-by-step. Whenever necessary, you may elaborate in more detail. Depending on the topic’s complexity, invest one to three paragraphs. The fewer, the better. If you find that there is insufficient information to verify the Claim, explicitly state what information is missing. If you cite web sources, always refer to them by including their URL as a Markdown hyperlink. Use information only from the recorded evidence: Avoid inserting information that is not implied by the evidence. You may use commonsense knowledge, though.
[Record]
Your Analysis:
E.5 Judge Prompt Template
Instructions Determine the Claim’s veracity by following these steps: 1. Briefly summarize the key insights from the fact-check (see Record) in at most one paragraph. 2. Write one paragraph about which one of the Decision Options applies best. Include the most appropriate decision option at the end and enclose it in backticks like ‘this‘.
[Extra Rules]
[Decision Options]
[Record]
Your Judgement:
@@ E.5 Judge Prompt Template
E.6 Justify Prompt Template
Instructions You are provided with the record of a fact-check. It contains the Claim to be verified and documentation of all the fact-checking work along with the gathered evidence. Your task is to summarize the fact-check. That is, you provide a concise, one-paragraph justification for the final VERDICT based on the knowledge from the Record. Note: Be truthful, brief, and do not add any additional information besides the information given in the Record. Link key sources in your summary. Use Markdown notation for that. You may link them in-line. Don’t state the Claim again. Rather focus on the key insights of the fact-check. Simply print just the summary.
[Record]
Summary:
Appendix F Wrongly Annotated VERITE Samples
@@
The wrong annotation can be seen even more clearly for the claims $249$ and $250$. DEFAME’s reverse search yielded multiple credible sources, including an article by CBS News1515footnotemark: 15, consistently reporting about a gathering of $75$ people at the “Area 51” military airbase in Sept 2019. 161616www.cbsnews.com/news/storm-area-51-hoax-draws-hundreds-events-outside-secretive-us-base-today-2019-09-20-live-updates/ The sources use the claim’s photo along with other similar images showing the apparently same event. According to the VERITE annotation, the photo shows a completely different event, contradicting the evidence. Consequently, the provided labels for both claims are clearly “switched.” Since we analyzed only $20$ samples, there are likely more such wrongly annotated samples, penalizing DEFAME’s accuracy where it should not be. Hence, the actual accuracy of DEFAME is slightly higher than measured.
@@
Appendix G Outdated Ground Truths and Temporal Dependence in MOCHEG
@@
Appendix H Details on the Human Evaluation
We ensure a minimum number of $5$ evaluations per claim and include random samples from the Chain-of-Thought baseline for comparison. (We post-process the baseline outputs to match the format of DEFAME’s output to “disguise” it.) In total, $154$ of the submissions assess the outputs from DEFAME while $31$ assess the baseline. An example of the report format is included in Appendix K. We assess the difference in Completeness scores between the DEFAME and CoT LLM groups using the Mann-Whitney U Test. This non-parametric test was chosen due to the non-normal distribution of completeness scores. The Mann-Whitney U Test yields a $p$-value of approximately $9.1\times 10^{-9}$, deeming the findings statistically significant. All Participants have a higher degree in education.
To provide further insights, we include a direct comparison of a fact-checking report generated by DEFAME and one from the Chain-of-Thought baseline for the same claim (see Figures  9 and  10). This example illustrates the key differences observed during the evaluation, particularly in the Completeness dimension. While both reports maintain coherence in structure and logical flow, the DEFAME report explicitly links its verdict to multiple pieces of evidence, providing a clear justification. In contrast, the CoT report relies heavily on parametric reasoning, lacking grounded evidence to support its conclusions.
@@ Appendix H Details on the Human Evaluation
Appendix I Automatic Justification Evaluation
@@
Appendix J ClaimReview2024+
To construct the ClaimReview2024+ dataset, we scraped 300 claims from the ClaimReview API provided by Google [^19] https://toolbox.google.com/factcheck/apis, covering the period from November 1, 2024, to January 18, 2024. The dataset consists of 129 [Uncaptioned image]Refuted, 89 [Uncaptioned image]Supported, 61 [Uncaptioned image]Misleading, and 21 [Uncaptioned image]NEI claims, with 160 being text-only and 140 paired with a corresponding image. While the labels follow the annotations provided by fact-checking websites, adjustments were made to align them with our taxonomy.
We applied a combination of automated and manual steps to standardize labeling. Automated classification was based on patterns in the original annotations, while ambiguous or conflicting cases were manually reviewed for consistency. Additionally, some claims were cleaned to remove explicit indications of veracity in the text. For example, a claim like ”Fact Check: Image from Pakistan falsely claims to show Punjab police using drugs” was reworded to ”Image from Pakistan shows Punjab police using drugs” to avoid embedding the claim’s outcome in its phrasing.
Finally, while this dataset was designed to contain claims novel to GPT-4o, we acknowledge the possibility that some claims may have appeared in earlier forms and were later re-evaluated by fact-checking websites before being picked up by ClaimReview.
Appendix K Example Fact-Checking Report
Figures 11 and 12 display a 2-page fact-checking report as returned by DEFAME, including a correct veracity prediction. We rendered the Markdown output into a PDF, including the referenced images. This and the following reports also include hyperlinks [^20] For technical reasons, the hyperlinks are not preserved in this PDF., referencing the used resources.
@@
Appendix L Examples of Failure Cases
The qualitative analysis of mispredicted VERITE instances uncovered two common failure modes attributed to DEFAME: premature judgment and failed evidence retrieval. See Figure 13 for an example of a premature judgment. While it is true that there exists a photo showing the “Air Force Officers’ Spouses’ Club” taken on April 24, 2018212121https://web.archive.org/web/20241116170848/https://taskandpurpose.com/military-life/viral-military-spouse-tweet-photo/, it is not the photo depicted in the claim. DEFAME missed comparing the claim photo with evidence photos before judging the veracity.
Additionally, Figures 14 and 15 show a case where the retrieval of evidence (from either (reverse) image search or web search) was unsuccessful, resulting in a wrong prediction. Manual inspection reveals that according to Snopes [^22] https://www.snopes.com/fact-check/sharks-on-power-lines-hurricane/, the origin is a news video accessible on YouTube. However, both Snopes and YouTube are excluded from DEFAME’s search results. Apart from that, manually using all three web search tools yields only more excluded or unrelated results.

### v3

@@ H.3 Summarize Prompt Template
H.4 Develop Prompt Template
Instructions You just retrieved new Evidence. Now, analyze the Claim’s veracity using the evidence. Always adhere to the following rules: Focus on developing new insights. Do not repeat larger parts from the Record. Do not restate the Claim. Write down your thoughts step-by-step. Whenever necessary, you may elaborate in more detail. Depending on the topic’s complexity, invest one to three paragraphs. The fewer, the better. If you find that there is insufficient information to verify the Claim, explicitly state what information is missing. If you cite web sources, always refer to them by including their URL as a Markdown hyperlink. Use information only from the recorded evidence: Avoid inserting information that is not implied by the evidence. You may use commonsense knowledge, though. [Record] Your Analysis:
H.5 Judge Prompt Template
Instructions Determine the Claim’s veracity by following these steps: 1. Briefly summarize the key insights from the fact-check (see Record) in at most one paragraph. 2. Write one paragraph about which one of the Decision Options applies best. Include the most appropriate decision option at the end and enclose it in backticks like ‘this‘. [Extra Rules] [Decision Options] [Record] Your Judgement:
@@ H.5 Judge Prompt Template
H.6 Justify Prompt Template
Instructions You are provided with the record of a fact-check. It contains the Claim to be verified and documentation of all the fact-checking work along with the gathered evidence. Your task is to summarize the fact-check. That is, you provide a concise, one-paragraph justification for the final VERDICT based on the knowledge from the Record. Note: Be truthful, brief, and do not add any additional information besides the information given in the Record. Link key sources in your summary. Use Markdown notation for that. You may link them in-line. Don’t state the Claim again. Rather focus on the key insights of the fact-check. Simply print just the summary. [Record] Summary:
Appendix I Wrongly Annotated VERITE Samples
@@
The wrong annotation can be seen even more clearly for the claims $249$ and $250$. DEFAME’s reverse search yielded multiple credible sources, including an article by CBS News1212footnotemark: 12, consistently reporting about a gathering of $75$ people at the “Area 51” military airbase in Sept 2019. 131313www.cbsnews.com/news/storm-area-51-hoax-draws-hundreds-events-outside-secretive-us-base-today-2019-09-20-live-updates/ The sources use the claim’s photo along with other similar images showing the apparently same event. According to the VERITE annotation, the photo shows a completely different event, contradicting the evidence. Consequently, the provided labels for both claims are clearly “switched.” Since we analyzed only $20$ samples, there are likely more such wrongly annotated samples, penalizing DEFAME’s accuracy where it should not be. Hence, the actual accuracy of DEFAME is slightly higher than measured.
@@
Appendix J MOCHEG Metric Clarification
Equivalence of Micro-F1 and Accuracy in Multi-Class Settings.
In MOCHEG’s single-label, multi-class setting, micro-F1 score and accuracy are mathematically equivalent.
To see this, recall that the micro-F1 score is defined as
[…]
Now observe that since each incorrect prediction contributes exactly one false positive and one false negative, and all correct predictions are true positives, we can write
TP+12(FP+FN)=Total # of $\text{TP}+\frac{1}{2}(\text{FP}+\text{FN})=\text{Total \# of predictions}.$
Therefore, the micro-F1 score becomes
[…]
This identity holds in any multiclass, single-label classification setting. We include this clarification to prevent confusion caused by legacy naming conventions in prior work on MOCHEG.
Appendix K Outdated Ground Truths and Temporal Dependence in MOCHEG
@@
Appendix L Details on the Human Evaluation
We ensure a minimum number of $5$ evaluations per claim and include random samples from the Chain-of-Thought baseline for comparison. (We post-process the baseline outputs to match the format of DEFAME’s output to “disguise” it.) In total, $154$ of the submissions assess the outputs from DEFAME while $31$ assess the baseline. An example of the report format is included in Appendix N. We assess the difference in Completeness scores between the DEFAME and CoT LLM groups using the Mann-Whitney U Test. This non-parametric test was chosen due to the non-normal distribution of completeness scores. The Mann-Whitney U Test yields a $p$-value of approximately $9.1\times 10^{-9}$, deeming the findings statistically significant. All Participants have a higher degree in education.
To provide further insights, we include a direct comparison of a fact-checking report generated by DEFAME and one from the Chain-of-Thought baseline for the same claim (see Figures  10 and  11). This example illustrates the key differences observed during the evaluation, particularly in the Completeness dimension. While both reports maintain coherence in structure and logical flow, the DEFAME report explicitly links its verdict to multiple pieces of evidence, providing a clear justification. In contrast, the CoT report relies heavily on parametric reasoning, lacking grounded evidence to support its conclusions.
@@
Appendix M Automatic Justification Evaluation
@@
Appendix N Example Fact-Checking Reports
Figures 12 to 15 display two fact-checking reports as returned by DEFAME, including a correct veracity prediction. We rendered the Markdown output into a PDF, including the referenced images. These and all following reports also include hyperlinks [^16] For technical reasons, the hyperlinks are not preserved in this PDF., referencing the used resources.
@@
Appendix O Examples of Failure Cases
The qualitative analysis of mispredicted VERITE instances uncovered two common failure modes attributed to DEFAME: premature judgment and failed evidence retrieval. See Figure 16 for an example of a premature judgment. While it is true that there exists a photo showing the “Air Force Officers’ Spouses’ Club” taken on April 24, 2018171717https://web.archive.org/web/20241116170848/https://taskandpurpose.com/military-life/viral-military-spouse-tweet-photo/, it is not the photo depicted in the claim. DEFAME missed comparing the claim photo with evidence photos before judging the veracity.
Additionally, Figures 17 and 18 show a case where the retrieval of evidence (from either (reverse) image search or web search) was unsuccessful, resulting in a wrong prediction. Manual inspection reveals that according to Snopes [^18] https://www.snopes.com/fact-check/sharks-on-power-lines-hurricane/, the origin is a news video accessible on YouTube. However, both Snopes and YouTube are excluded from DEFAME’s search results. Apart from that, manually using all three web search tools yields only more excluded or unrelated results.